Rohan was as old as his son is at present at the time of his birth.If rohan's age is 40 years now, then what is his son's age?
Answer by ptaylor(2198)

Let x=Rohan's son's age at present
40-x=Rohan's father's age at the time of his son's birth and this equals the son's present age
Sooooo
40-x=x
2x=40
x=20-----Rohan's son's present age
CK
40-20=20
Hope this helps---ptaylor